OpenStudy (anonymous):

Express the function h(x) = (1/(x+1)) in the form fog . If g(x) = x+1, find the function f(x). Your answer is f(x) _________, and with your choice for f(x) compute f(g(x)) = _____ = h(x) ?

OpenStudy (anonymous):

f(x)=1/x, because f(g(x))=h(x) given, and g(x)=(x+1) given, That means we have to plug the value of function g(x) in the function f(x) in place of x to get "fog".

OpenStudy (anonymous):

So f(x)=1/x and f(g(x))=1/(x+1)